An analyst has predicted the following returns for Stock A and Stock B in three possible states of the economy State Probabili Boom Normal Recession 0.25 .24 0.27 0.49 0.160.20 0.10 0.17 a. What is the probability of a recession? (Round your answer to 2 decimal places.) Probability 0.26 b. Calculate the expected return for Stock A and Stock B. (Round your answers to 2 decimal places Expected Return Stocks A Stocks B C. Calculate the expected return for a portfolio that is invested 42% in A and 58% in B. (Round your answer to 2 decimal pl aces.) Expected return
Answer #1

Solution : ( a )

Probability of a recession-1-(0.25 0.49) 1-0.74 = 0.26

========================================================================

Solution : ( b )

Expected returns for Stocks A = 0.25 * 0.24 0.49 * 0.16 + 0.26 * 0.10 = 0.06 0.0784 0.026 = 0.1644 16.44%

Expected returns for Stocks B = 0.25 * 0.27 0.49 * 0.20 0.26 * 0.17 0.0675 + 0.098 0.0442 = 0.2097 = 20.97%

========================================================================